250 tickets to a show were sold,some at 200 and others at R300 the total amount of money taken for the show was R57 500 how many R200 tickets and how many R300 tickets were sold
Answer by ikleyn(52787)

.
250 tickets to a show were sold,some at 200 and others at R300
the total amount of money taken for the show was R57 500
how many R200 tickets and how many R300 tickets were sold
~~~~~~~~~~~~~~~~

Let x be the number of R300 tickets.

Then the number of the R200 tickets is (250-x).


Write the revenue equation (the total money equation)

    300x + 200*(250-x) = 57500.


Simplify and find x

    300x + 200*250 - 200x = 57500

           100x           = 57500 - 200*250

           100x           =     7500

              x           =     7500/100 = 75.


ANSWER.  75 tickets at R300  and the rest  250-75 = 175 tickets at R200.


CHECK.   75*300 + 175*200 = 57500,  total money.   ! correct !

Solved.

------------

It is a standard and typical tickets problem.

There are different methods of solving such problems.
In this site, there is a lesson
    - Using systems of equations to solve problems on tickets
explaining and showing different methods of solving such problems.

From this lesson,  learn on how to solve such problems once and for all.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Systems of two linear equations in two unknowns".

Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.